OpenStudy (anonymous):

You drive on Interstate 10 from San Antonio to Houston, half the time at 55 km/h and the other half at 90 km/h. On the way back you travel half the distance at 55 km/h and the other half at 90 km/h. What is your average speed (a) from San Antonio to Houston, (b) from Houston back to San Antonio, and (c) for the entire trip? (d) What is your average velocity for the entire trip? (e) Sketch x versus t for (a), assuming the motion is all in the positive x direction. Indicate how the average velocity can be found on the sketch.

OpenStudy (anonymous):

(a)avg speed=(55+90)/2=145/2=72.5 km/h (b)Now from Houston To San Antonio,let total distance be 2d. Then total time taken=d/55 + d/90 Hence,avg speed=2d/{d/55 + d/90} =(2*55*90)/145=68.276 km/h